OpenStudy (anonymous):

can someone explain how to multiply 4 x 1 3/4 i forgot how to multiply fractions with whole numbers..

OpenStudy (anonymous):

\[4\times 1\tfrac{3}{4}=4\times \frac{7}{4}=\frac{28}{4}=7\]

OpenStudy (anonymous):

or if you prefer, \[4\times 1\tfrac{3}{4}=4\times 1+4\times \frac{3}{4}=4+3=7\] either way